OpenStudy (anonymous):

[(21 + 6) − 32] ÷ 9 ⋅ 2.

OpenStudy (texaschic101):

do parenthesis first, then brackets, then division, then multiplication

oregonduck (oregonduck):

so... first... 21+6=27 second... 27-32=-5

OpenStudy (anonymous):

what is the answer i cant seem to find it

oregonduck (oregonduck):

third... -5 divided by 9=? what is -5 divided by 9 @lukehutch100k

OpenStudy (texaschic101):

or just do -5/9 * 2

OpenStudy (anonymous):

0.5

oregonduck (oregonduck):

it is -0.55555555555 so -0.55*2= -1.1111...

oregonduck (oregonduck):

the answer to: [(21 + 6) − 32] ÷ 9 ⋅ 2= -1.111...

OpenStudy (anonymous):

thank youuu!

oregonduck (oregonduck):

np anytime m8